Abby consumes only apples. Year 1 Year 2 red $1 each $2 each green $2 each...

  1. Abby consumes only apples.

Year 1

Year 2

red

$1 each

$2 each

green

$2 each

$1 each

In year 1 Abby buys 10 red apples and in Year 2 Abby buys 10 green apples.

Compute a CPI for apples for each year. Assume that year 1 is the base year in which the consumer basket is fixed. How does your index change from year 1 to year 2? (at least five sentences to explain)

Suppose Abby is happy eating red or green apples. How much has the true cost of living increased for Abby? (at least five sentences to explain)
Answer #1

Year 1 is Base Year

Year 1: Price of Red Apples = $1 per apple

Price of Green Apples = $2 per apple

Year 2: Price of Red Apples = $2 per apple

Price of Green Apples = $1 per apple

Year 1: 10 red apples are bought

Cost of the Basket in current Period = $1 *10 = $10

Cost of the Basket in Base Period = $1 *10 = $10

CPI of Year 1 = (Cost of the Basket in current Period/Cost of the Basket in Base Period)*100=100

Year 2: 10 green apples are bought

Cost of the Basket in current Period = $1*10 = 10

Cost of the Basket in Base Period = $2 *10 = $20

CPI of Year 2 = (10/20)*100 = 50

The CPI change from 100 to 50 between Year 1 and Year 2

If Abby is happy eating either Red or Green apples, the cost of living has not changed. Thus, this is contrary to the CPI measure that shows that there has been decrease in the cost of living for Abby.